Carlos can strip a hardwood floor in 3 hours, agnew can strip the same area in 6 hours how long wuld it take the two workers working together?
Answer by stanbon(75887)

Carlos can strip a hardwood floor in 3 hours, agnew can strip the same area in 6 hours how long wuld it take the two workers working together?
---
Carlos rate: 1/3 job/hr
Agnew rate:: 1/6 job/hr
---
Together rate: 1/x job/hr
-----
Equation:
rate + rate = together rate
1/3 + 1/6 = 1/x
----
2x + x = 6
3x = 6
x = 2 hrs (time to do the job together)
=========================================
cheers,
Stan H.